# Not-so-glass-like Caging and Fluctuations of an Active Matter Model

### v1

@@ Abstract
The finite-time position of the first peak and its growth with dimension lead us to a direct geometrical interpretation. Figure 4 indeed reveals that the peak height first increases with $\hat{\varphi}$, saturates around $\hat{\varphi}_{d}$, and then decreases, and that $\widehat{\mathrm{Pe}}$ linearly enhances the effect. Given that $\hat{\alpha}_{2}$ encodes fluctuations in cage sizes, which are maximal around $\hat{\varphi}_{d}$ Biroli et al. (2022); Charbonneau et al. (2024), the first trend follows. It further highlights the critical origin of the phenomenon, with an actual divergence expected in the limit $d\rightarrow\infty$. The second trend is consistent with an active tracer exploring ever lower dimensional features – down to edges and wedges – of its surroundings upon approaching caging. The robustness and magnitude of this effect suggest its importance in capturing the dynamics of active glass formers. A proper theory of these systems would then be expected to capture this distinctive fluctuation-based signature.
@@
Figure 4: Density evolution of the height of the first peak of $\hat{\alpha}_{2}$ at $\widehat{\mathrm{Pe}}=10$ for various $d$ and obstacle densities in the ABP-RLG. The peak height generally saturates as $d$ increases, but is roughly proportional to $d$ around $\hat{\varphi}_{d}$ as in BP-RLG Charbonneau et al. (2024). (inset) The linear growth of the peak height with $\widehat{\mathrm{Pe}}$ for $d=12$ and $\hat{\varphi}=1$ (dashed line), $2.5$ (solid line) and $4$ (dot-dashed line).
@@
Studying the ABP-RLG provides various key insights into the physics of active glasses. In particular, the mean-field–like dynamical transition shifts to higher densities with activity and saturates at the percolation transition, thus identifying activity as a mechanism to sidestep what would be instantonic (or activated) processes in standard glass formers. Activity also leads to the exploration of lower-dimensional caging features, thus resulting in a clear and distinct signature in the NGP around $\widehat{\varphi}_{\mathrm{d}}$. This finite-dimensional echo of mean-field criticality is well within the reach of simulation and experiments in these systems and is therefore a signature with which to contrast standard and active glass formers. Furthermore, these findings highlight the key importance of solving the DMFT, as well as its small and large fluctuations, as the necessary path toward achieving a first-principle description of glasses, whether active or not.
@@
We would like to thank Ludovic Berthier, Giampaolo Folena, Yi Hu, and Grzegorz Szamel for many stimulating exchanges. This work was supported by a grant from the Simons Foundation (Grant No. 454937). Data relevant to this work have been archived and can be accessed at the Duke Digital Repository https://doi.org/10.7924/XXXXXXX.

### v2

@@ Abstract
The finite-time position of the first peak and its growth with dimension hint at a direct geometrical interpretation. Figure 4 reveals that the peak height first increases with $\hat{\varphi}$, saturates around $\hat{\varphi}_{d}$, and then decreases. It also shows that $\widehat{\mathrm{Pe}}$ linearly enhances the effect (inset) SI . Because $\hat{\alpha}_{2}$ encodes fluctuations in cage sizes, which are maximal around $\hat{\varphi}_{d}$ Biroli et al. (2022); Charbonneau et al. (2024), the explanation for the peak position naturally follows. It further highlights the critical origin of the phenomenon, with an actual divergence expected in the limit $d\rightarrow\infty$. The enhancement of the effect with $\widehat{\mathrm{Pe}}$ is consistent with an active tracer exploring ever lower dimensional features – down to edges and cavities formed by obstacles – of its surroundings upon approaching caging. The robustness and magnitude of this effect suggest its importance in capturing the dynamics of active glass formers. A proper theory of these systems would then be expected to capture this distinctive fluctuation-based signature.
@@
Figure 4: Density evolution of the height of the first peak of $\hat{\alpha}_{2}$ for the ABP-RLG at $\widehat{\mathrm{Pe}}=10$ for various $d$ and obstacle densities. The peak height generally saturates as $d$ increases, but is roughly proportional to $d$ around $\hat{\varphi}_{d}$ similarly to the long-time peak of the BP-RLG Charbonneau et al. (2024). (inset) The linear growth of the peak height with $\widehat{\mathrm{Pe}}$ for $d=12$ and $\hat{\varphi}=1$ (dashed line), $2.5$ (solid line) and 44 (dot-dashed line).
@@
The ABP-RLG provides key insights into the physics of dense active systems. First, its mean-field–like dynamical transition shifts to higher densities with activity and saturates at the percolation transition, thus identifying activity as a mechanism that sidesteps the instantonlike (or activated) hopping processes of standard glass formers. Second, an active tracer tends to explore lower-dimensional caging features, resulting in a clear and distinct short-time peak of the NGP around $\widehat{\varphi}_{\mathrm{d}}$. This finite-dimensional echo of mean-field criticality is well within the reach of simulation and experiments in these systems and is therefore a distinctive feature with which to contrast standard and active glass formers. These findings could further be used to design systems in which the efficacy of active particle caging can be tuned, especially for $d>2$. Finally, these findings once again highlight the importance of solving the DMFT, as well as its small and large fluctuation corrections, as the necessary path toward achieving a first-principle description of glasses, both active and not.
Acknowledgements.
@@
We thank Ludovic Berthier, Giampaolo Folena, Yi Hu, Grzegorz Szamel, and Yikang Zhang for many stimulating exchanges. This work was supported by a grant from the Simons Foundation (Grant No. 454937). D. K. thanks the European Union Horizon 2020 research and innovation program under Grant Agreement No. 857470 and from the European Regional Development Fund via the Foundation for Polish Science International Research Agenda PLUS program GrantNo. MAB PLUS/2018/8, and the Ministry of Science and Higher Education’s initiative “Support for the Activities of Centers of Excellence Established in Poland under the Horizon 2020 Program” (agreement no. MEiN/2023/DIR/3795
Data availability –
Data relevant to this work have been archived and can be accessed at the Duke Research Data Repository Zheng et al. (2025).